### v1

# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

Abstract
In this paper, we showcase how flow obstruction by a deformable object can lead to symmetry breaking in curved domains subject to angular acceleration. Our analysis is motivated by the deflection of the cupula, a soft tissue located in the inner ear that is used to perceive rotational motion as part of the vestibular system. The cupula is understood to block the rotation-induced flow in a toroidal region with the flow-induced deformation of the cupula used by the brain to infer motion. By asymptotically solving the governing equations for this flow, we characterise regimes for which the sensory system is sensitive to either angular velocity or angular acceleration. Moreover, we show the fluid flow is not symmetric in the latter case. Finally, we extend our analysis of symmetry breaking to understand the formation of vortical flow in cavernous regions within channels. We discuss the implications of our results for the sensing of rotation by mammals.
keywords:
1 Introduction
Rotational motion of the head in humans is perceived through the vestibular system, which is located in the inner ear (Paulin & Hoffman, 2019). For mathematical modeling purposes, this system can be described as a set of three mutually orthogonal, nearly circular canals, known as the semicircular canals (SCCs) in the anatomical literature (Oghalai & Brownell, 2020; Curthoys & Oman, 1987). These canals resemble deformed tori, where the slender regions are filled with a Newtonian fluid called endolymph. The larger region, termed the utricle, houses a gelatinous protein-polysaccharide elastic membrane known as the cupula (Casale et al., 2024), which is innervated by hair cells (cilia); the innervated cilia transmit mechanical deflections of the cupula to the nervous system via the vestibular nerve (Waxman, 2024). In mechanical terms, a change in angular velocity about a given axis drives a fluid flow in that canal, which generates a pressure gradient, deforming the cupula (and hence the innervated hairs) and thus allowing the brain to sense the motion. Specifically, as the walls of the canal rigidly rotate with the head, the fluid in the center is left behind, causing the cupula to deform in the opposite direction to the imposed rotation. The mutually orthogonal structure of the semicircular canals allows the detection of any three dimensional rotation of the head. A schematic diagram of the vestibular system is provided in Figure 1.
Figure 1: Schematic of the vestibular system. (a) The inner ear and vestibular apparatus: Three mutually orthogonal semicircular canals (SCCs), each containing a cupula, send information to the nervous system about the rotational motion of the head. (b) Zoom in of the obstruction within each SCC caused by the cupula. Information about the rotation of each SCC is inferred from the deflection of its cupula — the inertia of the fluid that fills the SCC (endolymph) causes the cupula to deform. (Cupula deformation is sensed via innervated cilia that are embedded within the cupula.)
The vestibular system can be affected by several pathologies that disrupt its normal function. One of the most common disorders is benign paroxysmal positional vertigo (BPPV), in which brief episodes of vertigo are triggered by specific head movements. BPPV occurs when calcium carbonate crystals (otoconia) dislodge from the utricle and move into the semicircular canals, causing abnormal stimulation of the vestibular nerve (Hornibrook, 2011). Another significant condition is vestibular neuritis, an inflammation of the vestibular nerve, usually caused by viral infections, which leads to acute vertigo, imbalance and nausea (Royal & Vargas, 2014). Meniere’s disease also affects the vestibular system, causing episodic vertigo due to abnormal fluid buildup in the inner ear, leading to disturbances in balance (Harcourt et al., 2014). Early diagnosis and appropriate treatment of these vestibular pathologies are essential to improving quality of life and preventing chronic balance issues. Here, mathematical modelling has great potential in enabling for quantitative predictions of balance response and in elucidating the sensitivity of the vestibular system to material changes, for instance as may occur with ageing (Konrad et al., 1999).
A number of mathematical models exist for the vestibular dynamics, both numerical and analytical. On the analytical side, previous models can be classified into two broad categories. The first approach is that of Obrist and co-authors (Obrist, 2008; Grieser et al., 2012; Vega et al., 2008) in which the geometry is idealized to allow a solution to be found under arbitrary forcing of the system, i.e. arbitrary rotational motion. The second approach is that of Rabbitt & Damiano, who maintain a more realistic geometry but require strong assumptions on the form of the forcing to make analytical progress. In particular, in their series of papers (Rabbitt & Damiano, 1992; Damiano & Rabbitt, 1996; Damiano, 1999), Rabbitt & Damiano assume that the forcing is sinusoidal, as might be expected when moving the head up or down (for instance when nodding). Fully numerical investigations of the vestibular system also exist (Chung, 2010; Goyens et al., 2019; Boselli et al., 2009, 2013) — these generally implement a realistic channel geometry but do not model the cupular deformation as a fluid-structure interaction. Instead, they incorporate the effect of cupular deformation via a periodic boundary condition for the flow coupled to a time dependent pressure jump.
In this paper, we present both numerical simulations and an analytical approach for cupular dynamics. Our analytical model is derived from first principles, including explicitly both toroidal fluid flow and the mechanics of the cupula. By exploiting the slenderness of the semicircular canals, and applying a detailed asymptotic analysis, we obtain a reduced model that allows to incorporate both arbitrary geometry and arbitrary forcing, combining the best of previous approaches. We complement this with numerical computations, specifically including fluid-solid couplings. Our combined numerical and analytical approach enables us to validate the reduced analytical model, uncover qualitative features, characterize flow regimes, and investigate detailed features of the flow and cupula deformation.
One of the key motivating issues underlying our study concerns the mechanical properties of the cupula. Although the anatomy of the vestibular system is well understood, the architecture itself is incredibly delicate and fragile, which prohibits the possibility of direct mechanical testing. For this reason, the stiffness of the cupula has only ever been obtained through indirect measurement, a procedure that has produced both some uncertainty and surprisingly low values; for example, a Young’s modulus of around 5 Pa has been reported (Selva et al., 2009), which is well below values typically associated with soft biological tissues (see Budday et al., 2015, where they estimate the Young’s modulus of brain matter to be 1 kPa). Nevertheless, the stiffness of the cupula is a key mechanical parameter, as it dictates the degree of deformation under a given flow and therefore the potential for and degree of stimulus. In fact, as we will show, this parameter plays an even stronger role, impacting not just the degree of deformation but the qualitative nature of the flow induced by motion as well. By examining the behaviour of our model as the relative stiffness of the cupula varies, we will demonstrate the presence of two distinct regimes: for “soft” cupulas the deformation follows the angular velocity of the imposed motion, while for “stiff” cupulas the deformation instead tracks the angular acceleration.
Moreover, we will explain how the second of these regimes is connected to a symmetry breaking of the flow in the endolymph. As we shall demonstrate in our numerical simulations, presented in Section 2.2, the flow in the endolymph is only axisymmetric (relative to the duct’s center-line or axis) under particular conditions. Despite this observation, many existing models have implicitly assumed a symmetric flow. Examination of our analytical solution enables us to explain exactly when and how symmetry breaking occurs. This feature is interesting in the context of the broader literature on flow through curved pipes. Although the effect of curvature on pipe flow was first discussed by Dean (1928), the plethora of subsequent studies have focused mostly on steady flows (Chupin & Stepanov, 2008; Pedley, 1980; Siggers & Waters, 2005) — there have been many fewer investigations into unsteady fluid phenomena (though see Siggers & Waters, 2008, for an example). We shall show that the essential coupling between the Euler force and the (a priori) unknown pressure gradient can lead to the annihilation of the symmetric leading order velocity — a situation that distinguishes this problem from classical studies of flow in curved pipes. We conclude our study with an investigation of the emergence of vortical flow in the utricle. This feature has been reported previously, but only in numerical simulations (Boselli et al., 2013); our model provides both an analytical understanding and an explicit characterization for when vortical flow will emerge.
2 Governing equations
Figure 2: Problem setup. (a) Plan view of a semicircular canal showing the spatially-varying canal radius, $a(s)$, and the cupula (shaded in grey), which is situated in the enlarged portion, or utricle. (b) Schematic of the chosen coordinate system. (c) Close up of the region around the cupula, highlighting the cupula’s thickness, $t_{h}$, and its attachment to the canal walls via the ‘crista’ (black region). (The toroidal flow is shown schematically here to allow the zoom in on the cupula.)
We consider a single semicircular canal, as portrayed schematically in Figure 2(a): the endolymph fills a toroidal structure whose centreline forms a circle of radius $R$ and whose radius is small and spatially-varying, denoted $\hat{a}(\hat{s})\ll R$, where $\hat{s}$ is an arc length parameter along the centreline. The canal is subjected to a rotation defined by angular velocity $\hat{\Omega}(\hat{t})$ around the centre of the toroid with rotation axis normal to the plane of the centreline. The endolymph is assumed to be an incompressible Newtonian fluid of dynamic viscosity $\mu$ and density $\rho$. The elastic, gel-like cupula occupies a thin region (shown in grey in fig. 2) and has density $\rho_{s}$, Young’s modulus $E$, thickness $t_{h}$ and Poisson ratio $\nu_{s}$. The cupula is assumed to occupy the entire cross-section of the canal, as can be seen in Figure 2(c), where the solid cupula is shaded in yellow and the liquid endolymph is shaded in blue. (The region shaded in black represents a structure called the crista which attaches the cupula to the canal wall. However, the azimuthal size of the crista will not appear directly in this work.)
2.1 Equations for the bulk
The Navier–Stokes equations for the dimensional fluid velocity $\hat{\mathbf{u}}$ and pressure $\hat{p}$ in the co-rotating frame are given by (Landau & Lifshitz, 1987)
$\boldsymbol{\nabla}\cdot\hat{\mathbf{u}}$ $=0,$ (1a)
$\frac{D\mathbf{\hat{u}}}{D\hat{t}}+\frac{\partial\mathbf{\hat{\Omega}}}{\partial\hat{t}}\times\mathbf{\hat{x}}+2\mathbf{\hat{\Omega}}\times\mathbf{\hat{u}}+\mathbf{\hat{\Omega}}\times(\mathbf{\hat{\Omega}}\times\mathbf{\hat{x}})$ $=-\boldsymbol{\nabla}\hat{p}+\mu\boldsymbol{\nabla}^{2}\mathbf{\hat{u}}.$ (1b)
Here, the first of the extra terms on the left hand side is the Euler force, the second term is the Coriolis force and the final additional term is the centrifugal force, each due to the imposed rotation. The fluid is assumed to satisfy the no-slip condition at the edges of the walls, so that $\mathbf{\hat{u}}(\hat{r}=\hat{a}(\hat{s}))=\mathbf{0}$ for $s\in\,(0,2\pi R)$. Motivated by the small strains in the cupula (Selva et al., 2009), it is modeled as a linearly elastic material, satisfying the steady Navier equations:
$\boldsymbol{\nabla}\cdot\boldsymbol{\hat{\tau}}=\mathbf{0},$ (2)
alongside a linear Hookean constitutive law relating stress $\boldsymbol{\hat{\tau}}$ and the strains. Finally, we model the fluid structure interaction at the cupula-endolymph boundary in the usual way, imposing continuity of velocity and stress (Gkanis & Kumar, 2006).
2.2 Numerical simulations
The system of equations (1)-(2) was simulated in COMSOL for different values of the Young’s modulus of the cupula and with a Poisson ratio $\nu_{s}=0.3$. We impose a simple sinusoidal forcing, given by $\hat{\Omega}(\hat{t})=\Omega_{0}\sin(2\pi t/t_{\rm typ})$, with $\Omega_{0}=1\mathrm{rad/s}$ and $t_{\rm typ}=1\mathrm{~{}s}$. The geometrical parameters are $a=1.6\times 10^{-4}$ m, $R=3.2\times 10^{-3}$ m and $t_{h}=0.8\times 10^{-4}$ m (Daocai et al., 2014).
The equations were solved for a solid cupula of finite size, i.e. with no thin cupula assumption. Further details on the numerical simulations are available in Appendix A. This includes deformation profiles of the solid cupula (Figure 12). Moreover, Figure 12 shows that the magnitude of cupula deformation is inversely proportional to Young’s modulus.
Flow profiles on either side of the cupula produced by these numerical simulations are shown in fig. 3. Here a top view of the mid-plane of the flow around the canal is plotted, with colour indicating the speed of the flow, normalized by the maximum speed throughout the flow and with fast regions coloured red, stagnant regions coloured blue.
The cupula appears as the central region and is shown in its deformed configuration. This deflection is imperceptible for all except the $E=10^{2}\mathrm{~{}Pa}$ case and so the relative magnitude of the cupula deformation is indicated by the grey scale colouring, which shows that it is maximum at the centre.
In the panels of fig. 3, the Young’s modulus, $E$, of the cupula increases from left to right. As should be expected, the magnitude of the deflection decreases as the material becomes stiffer. This is confirmed quantitatively in Figure 12 of Appendix A. Surprisingly, however, we also observe a significant change in flow behaviour: for small values of $E$, the flow is axially symmetric about the centerline of the tube but as $E$ increases the flow transitions, losing its axisymmetry and exhibiting vortical structures. It is worth noting that these plots use values of the Young’s modulus that cover a typical physiological range for soft biological tissues (Goriely, 2017). Notwithstanding our remark in the introduction that even lower values have been suggested via indirect methods, a flow transition with physiologically relevant values of $E$ suggests that asymmetric flow may be present in a physiological vestibular system. If so, this would contradict the assumption of axial symmetry that is typical in previous models (Obrist, 2008; Rabbitt & Damiano, 1992), and raises interesting questions about what impact such asymmetry might have on the mechanics of balance and rotational sensing. To study this behaviour further, we thus turn now to a theoretical analysis of the governing system.
Figure 3: Velocity fields in the cupula as computed using COMSOL simulations. Results are shown for a range of cupula stiffnesses. Colour represents the relative magnitude of the fluid speed, with red denoting regions in which the flow is fast and blue representing stagnant regions; streamlines are represented by solid black curves. As the stiffness of the cupula increases, a symmetry-breaking of the flow occurs. In particular, for values of the Young’s modulus $E>10^{3}$ the flow is usually not axially-symmetric. Here $\hat{\Omega}(\hat{t})=\sin\left(2\pi\hat{t}/\mathcal{T}\right)$ and the snapshots are taken at $\hat{t}=0.25$ s and the period is $\mathcal{T}=1$ sec.
2.3 Theoretical model
To investigate the symmetry breaking observed in the numerical simulations, and to obtain a qualitative understanding of flow and deformation characteristics, in this section we use asymptotic analysis to derive a reduced order equation for the deformation of the cupula.
To capture the geometry of the semicircular canal, we introduce toroidal coordinates $(\hat{r},\theta,\hat{s})$, in which $\hat{s}\,\in\,(0,2\pi R)$ is the arc-length along the centreline of the torus. A sketch of this coordinate system is provided in Figure 2(b). Cartesian coordinates are related to toroidal coordinates via (Pedley, 1980):
$\begin{cases}\hat{x}=&(R+\hat{r}\cos{\theta})\cos\frac{\hat{s}}{R},\\ \hat{y}=&(R+\hat{r}\cos\theta)\sin\frac{\hat{s}}{R},\\ \hat{z}=&-\hat{r}\sin\theta.\end{cases}$ (3)
The negative sign in the last equation ensures that the orthonormal basis vectors $\{\mathbf{e}_{r},\mathbf{e}_{\theta},\mathbf{e}_{s}\}$ follow the right hand rule. We can now rewrite the Navier-Stokes equations (1) in component form. Writing the velocity vector as $\mathbf{\hat{u}}=\hat{u}\mathbf{e}_{r}+\hat{v}\mathbf{e}_{\theta}+\hat{w}\mathbf{e}_{s}$, the continuity equation becomes (see Pedley, 1980, for example)
$\frac{\partial\hat{u}}{\partial\hat{r}}+\frac{\hat{u}}{\hat{r}}+\frac{1}{\hat{r}}\frac{\partial\hat{v}}{\partial\theta}+\frac{1}{h}\frac{\partial\hat{w}}{\partial\hat{s}}-\frac{\hat{v}\sin\theta}{Rh}+\frac{\hat{u}\cos\theta}{Rh}=0,$ (4)
where $h=1+\hat{r}\cos(\theta)/R$ is a scale factor. The momentum equations are (see Pedley, 1980, for the equations expressed in an inertial frame):
$\begin{split}\rho&\left(\frac{\partial\hat{u}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{u}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\hat{v}^{2}}{\hat{r}}-\frac{\hat{w}^{2}}{h}\frac{\cos\theta}{R}-\hat{\Omega}\hat{w}\cos\theta-\hat{\Omega}^{2}Rh\cos\theta\right)\\ &=-\frac{\partial\hat{p}}{\partial\hat{r}}+\frac{\mu}{\hat{r}h}\left[\frac{\hat{r}}{h}\frac{\partial}{\partial\hat{s}}\left(\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\partial}{\partial\hat{r}}(h\hat{w})\right)-\frac{1}{\hat{r}}\frac{\partial}{\partial\theta}\left(\frac{h}{\hat{r}}\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{h}{\hat{r}}\frac{\partial\hat{u}}{\partial\theta}\right)\right],\end{split}$ (5a)
$\begin{split}\rho&\left(\frac{\partial\hat{v}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{v}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{v}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{v}}{\partial\hat{s}}+\frac{\hat{u}\hat{v}}{\hat{r}}+\frac{\hat{w}^{2}}{Rh}\sin\theta+\hat{\Omega}\hat{w}\sin\theta+\hat{\Omega}^{2}Rh\sin\theta\right)\\ &=-\frac{1}{\hat{r}}\frac{\partial\hat{p}}{\partial\theta}+\frac{\mu}{h}\frac{\partial}{\partial\hat{r}}\left[\frac{h}{\hat{r}}\left(\frac{\partial}{\partial\hat{r}}(\hat{r}\hat{v})-\frac{\partial\hat{u}}{\partial\theta}\right)\right]-\frac{\mu}{\hat{r}h^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial\hat{w}}{\partial\hat{s}}\right)-\hat{r}\frac{\partial^{2}\hat{v}}{\partial\hat{s}^{2}}\right],\end{split}$ (5b)
$\begin{split}\rho&\left(\frac{\partial\hat{w}}{\partial\hat{t}}+\hat{u}\frac{\partial\hat{w}}{\partial\hat{r}}+\frac{\hat{v}}{\hat{r}}\frac{\partial\hat{w}}{\partial\theta}+\frac{\hat{w}}{h}\frac{\partial\hat{w}}{\partial\hat{s}}+\frac{\hat{u}\hat{w}}{Rh}\cos\theta-\frac{\hat{v}\hat{w}}{Rh}\sin\theta\right.\\ &\left.+\frac{d\hat{\Omega}}{d\hat{t}}(R+\hat{r}\cos\theta)+\hat{\Omega}\hat{u}\cos\theta-\hat{\Omega}\hat{v}\sin\theta\right)\\ &=-\frac{1}{h}\frac{\partial\hat{p}}{\partial\hat{s}}+\frac{\mu}{\hat{r}^{2}}\frac{\partial}{\partial\theta}\left[\frac{1}{h}\frac{\partial}{\partial\theta}(h\hat{w})-\frac{\hat{r}}{h}\frac{\partial\hat{v}}{\partial\hat{s}}\right]-\frac{\mu}{\hat{r}}\frac{\partial}{\partial\hat{r}}\left[\frac{\hat{r}}{h}\left(\frac{\partial\hat{u}}{\partial\hat{s}}-\frac{\partial}{\partial\hat{r}}(h\hat{w})\right)\right].\end{split}$ (5c)
Noting that the cupula is approximately thin (with an aspect ratio of around 0.25), we model it as a circular plate. We locate the cupula at arc length position $\hat{s}=0$, and define its deflection by the function $\hat{\eta}(\hat{r},\theta,\hat{t})$. As indicated above, our numerical simulations have shown that the deformation of the cupula is small compared to the tube radius, suggesting that strains are small and suggesting it is sufficient to use a linear equation. This small strain assumption will be confirmed in Section 2.4 through a scaling argument. Motivated by this linearity and the relatively slender geometry of the cupula, we further assume that the cupular deflection satisfies the linear plate equation, which in the rotating frame reads (see Howell et al., 2008)
$B\nabla^{4}\hat{\eta}-T\nabla^{2}\hat{\eta}+\rho_{s}t_{h}\left(\frac{\partial^{2}\hat{\eta}}{\partial\hat{t}^{2}}+Rh\frac{\mathrm{d}\hat{\Omega}}{\mathrm{d}\hat{t}}\right)=\Delta\hat{p}=\hat{p}(\hat{r},\theta,\hat{s}=2\pi R,\hat{t})-\hat{p}(\hat{r},\theta,\hat{s}=0,\hat{t}).$ (6)
where $B=Et_{h}^{3}/[12(1-\nu_{s}^{2})]$ is the plate’s bending stiffness and $T$ is the applied tension in the plate (which we assume is uniform and constant since deflections remain small compared to $t_{h}$). In this regime, the forcing for the cupular displacement is given by the pressure jump across the two sides of the cupula.
We require boundary conditions for both the fluid problem (5) and the elastic problem (6). The walls of the endolymph are assumed to have a no-slip condition, so that in the rotating frame, $\mathbf{\hat{u}}(\hat{r}=\hat{a})=0$. We require a second set of boundary conditions where the cupula and the endolymph meet. The kinematic boundary condition requires that the cupular velocity and the endolymph velocity at the surface of the cupula must match. As the spatial gradients of the cupula’s deformation are small, we can write this as
$\frac{\partial\hat{\eta}}{\partial\hat{t}}\boldsymbol{e}_{s}=\mathbf{\hat{u}}(\hat{r},\hat{s}=0,\hat{t})=\mathbf{\hat{u}}(\hat{r},\hat{s}=2\pi R,\hat{t}).$ (7)
We also require boundary conditions for the plate equation (6). The precise attachment between the cupula and the utricle is an open area of research and the conditions to be satisfied are not immediately clear. We opt to implement a straightforward choice suggested by Rabbitt & Damiano (1992), which is to apply clamped boundary conditions at the edge of the cupula, i.e.
$\hat{\eta}(\hat{r}=\hat{a}(0))=\left.\frac{\partial\hat{\eta}}{\partial\hat{r}}\right|_{\hat{r}=\hat{a}(0)}=0.$ (8)
We also require initial conditions for the velocity, $\mathbf{\hat{u}}(\mathbf{\hat{x}},\hat{t}=0)$ and the cupular deflection $\hat{\eta}(\hat{r},\theta,\hat{t}=0)$. Unless otherwise stated we assume that the system is initially at rest and is undeformed.
2.4 Scalings and non-dimensionalization
The SCCs in mammals are thin and slender: with an aspect ratio $\epsilon=a/R$ between 0.05 and 0.1 (Daocai et al., 2014). It is therefore natural to exploit $\epsilon\ll 1$ and use an asymptotic approach to perform a long wavelength asymptotic analysis similar to lubrication theory. We also know from the continuity equation that if $\hat{w}\sim\mathcal{U}$ then $\hat{u},\hat{v}\sim\epsilon\mathcal{U}$. This is required to have a non-trivial balance and is a typical scaling in lubrication theory (Craster & Matar, 2006; Papageorgiou, 1995). To make this asymptotic intuition more formal, we non-dimensionalize, scaling the dimensional variables according to:
$\hat{r}=ar,\quad\hat{s}=Rs,\quad\hat{t}=\mathcal{T}t,\quad\hat{w}=\frac{R\dot{\Omega}_{0}a^{2}}{\nu}w,$ (9)
$\hat{p}=\frac{BR\Omega_{0}}{a^{2}\nu}p,\quad\hat{\eta}=\frac{a^{2}R\Omega_{0}}{\nu}\eta,\quad\frac{d\hat{\Omega}}{d\hat{t}}=\dot{\Omega}_{0}f(t),\quad\hat{\Omega}=\Omega_{0}F(t).$ (10)
(‘Unhatted’ variables are therefore dimensionless counterparts of the corresponding hatted variables.) Here $\mathcal{T}$ is the timescale of variation of the forcing so that $\dot{\Omega}_{0}=\Omega_{0}/\mathcal{T}$. The velocity scale $\mathcal{U}=R\dot{\Omega}_{0}a^{2}/\nu$ is chosen to balance the viscous forces with the Euler force. The pressure scale is chosen to balance the bending term and the forcing term in the plate equation (6). Note that $F(t)$ is the dimensionless angular velocity and $f(t)$ is the dimensionless angular acceleration so that by definition $f(t)=\dot{F}(t)$.
2.4.1 Dimensionless equations
Following the rescaling of equations (4)–(7), the dimensionless continuity equation (4) reads
$\frac{\partial u}{\partial r}+\frac{u}{r}+\frac{1}{r}\frac{\partial v}{\partial\theta}+\frac{1}{h}\frac{\partial w}{\partial s}-\epsilon\frac{v\sin\theta}{h}+\epsilon\frac{u\cos\theta}{h}=0,$ (11)
where the scale factor $h$ can be expressed as $h=1+\epsilon r\cos\theta$. The dimensionless Navier–Stokes equations (5) along the $\mathbf{e}_{r}$, $\mathbf{e}_{\theta}$ and $\mathbf{e}_{s}$ directions are given, respectively, by:
$\begin{split}\epsilon&\mathrm{St}\frac{\partial u}{\partial t}+\epsilon^{3}\mathrm{Re}\left(u\frac{\partial u}{\partial r}+\frac{v}{r}\frac{\partial u}{\partial\theta}+\frac{w}{h}\frac{\partial u}{\partial s}-\frac{v^{2}}{r}-\frac{1}{\epsilon}\frac{w^{2}}{h}\cos\theta\right)\\ &-\dot{\Omega}_{0}\mathcal{T}^{2}\mathrm{St}F(t)w\cos\theta-\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}h\cos\theta\\ &=-\kappa\frac{1}{\epsilon}\frac{\partial p}{\partial r}+\frac{\epsilon}{rh}\left[\frac{r}{h}\frac{\partial}{\partial s}\left(\epsilon^{2}\frac{\partial u}{\partial s}-\frac{\partial}{\partial r}(hw)\right)-\frac{1}{r}\frac{\partial}{\partial\theta}\left(\frac{h}{r}\frac{\partial}{\partial r}(r\hat{v})-\frac{h}{r}\frac{\partial\hat{u}}{\partial\theta}\right)\right],\end{split}$ (12a)
$\begin{split}\epsilon&\operatorname{St}\frac{\partial v}{\partial t}+\epsilon^{2}\operatorname{Re}\left(\epsilon u\frac{\partial v}{\partial r}+\epsilon\frac{v}{r}\frac{\partial v}{\partial\theta}+\epsilon\frac{w}{h}\frac{\partial v}{\partial s}+\epsilon\frac{uv}{r}+\frac{w^{2}}{h}\sin\theta\right)\\ &+\dot{\Omega}_{0}\mathcal{T}^{2}\mathrm{St}F(t)w\sin\theta+\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}h\sin\theta\\ &=-\kappa\frac{1}{\epsilon}\frac{1}{r}\frac{\partial p}{\partial\theta}+\frac{\epsilon}{h}\frac{\partial}{\partial r}\left[\frac{h}{r}\left(\frac{\partial}{\partial r}(rv)-\frac{\partial u}{\partial\theta}\right)\right]-\frac{\epsilon}{rh^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial w}{\partial s}\right)-\epsilon^{2}r\frac{\partial^{2}v}{\partial s^{2}}\right],\end{split}$ (12b)
$\begin{split}\mathrm{St}&\frac{\partial w}{\partial t}+f(t)(1+\epsilon r\cos\theta)+\epsilon\mathrm{St}\mathcal{T}^{2}\dot{\Omega}_{0}F(t)(u\cos\theta-v\sin\theta)\\ +&\epsilon^{2}\operatorname{Re}\left(u\frac{\partial w}{\partial r}+\frac{v}{r}\frac{\partial w}{\partial\theta}+\frac{w}{h}\frac{\partial w}{\partial s}+\epsilon\frac{uw}{h}\cos\theta-\epsilon\frac{vw}{h}\sin\theta\right)\\ &=-\kappa\frac{1}{h}\frac{\partial p}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left[\frac{r}{h}\left(\frac{\partial}{\partial r}(hw)-\epsilon^{2}\frac{\partial u}{\partial s}\right)\right]+\frac{1}{r^{2}}\frac{\partial}{\partial\theta}\left[\frac{1}{h}\frac{\partial}{\partial\theta}(hw)-\epsilon^{2}\frac{r}{h}\frac{\partial v}{\partial s}\right].\end{split}$ (12c)
The dimensionless form of (6) is given by
$\nabla^{4}\eta-b^{-1}\nabla^{2}\eta+\varrho\left(\operatorname{St}\frac{\partial^{2}\eta}{\partial t^{2}}+\frac{d\Omega}{dt}\right)=\Delta p.$ (13)
The non-dimensionalization procedure introduces several dimensionless parameters. We shall see that the most important of these are the stiffness
$\kappa=\frac{B\mathcal{T}}{a^{2}R\mu},$ (14)
which measures the time scale of forcing to the time scale of the cupula’s relaxation, and the Stokes number
$\mathrm{St}=\frac{a^{2}}{\nu\mathcal{T}},$ (15)
which measures the time scale of vorticity diffusion across the channel width, $a^{2}/\nu$, to the time scale of motion, $\mathcal{T}$. (Note that this version of the Stokes number arises in Stokes’ second problem and is sometimes replaced by the Womersley number, $\mathrm{Wm}=\mathrm{St}^{1/2}$.) We also introduce the Reynolds number of the flow, $\mathrm{Re}=\rho\mathcal{U}R/\mu$, as well as the cupula’s inertia, $\varrho=\rho_{s}a^{2}t_{h}\nu/(B\mathcal{T})$ and inverse bendability $b=B/(Ta^{2})$ (Davidovitch et al., 2011). To understand the relative size (and hence importance) of these parameters, we now discuss characteristic parameter values and typical sizes of dimensionless parameters next.
2.4.2 Parameter values
First, we consider the geometrical parameters, taken from Daocai et al. (2014): $R\approx 3.2\times 10^{-3}$ m, and $a\approx 1.6\times 10^{-4}$ m, so that the aspect ratio is $\epsilon\sim 0.05$. The cupula’s thickness is usually quoted as $t_{h}\approx 400\,\mu$m.
[…]
The most challenging parameters to identify are $B$ and $T$. We can infer the value of the bending stiffness from the results of Selva et al. (2009), who suggest an extremely low value of the Young’s modulus, $E\sim 5$ Pa. Other authors quote simply an estimate for the bending stiffness, $B\sim 10^{-10}\mathrm{~{}N\,m}$ (Rabbitt & Damiano, 1992). Since the cupula is difficult to access and measure, determining the pretension $T$ is significantly more challenging and typically requires detailed experiments. However, incorporating $T>0$ adds no additional complexity to our method of solution, and so we retain it for completeness; the effect of this tension is characterized through the inverse bendability parameter, $b=B/(Ta^{2})$ Davidovitch et al. (2011). We refer to the limit $b\to\infty$ as the plate regime and $b\to 0$ as the membrane regime.
In terms of the motion, usual ranges of operation for humans are $\dot{\Omega}_{0}\sim 1$ s-2 and $\mathcal{T}$ can be anywhere between 0.01 and 10 seconds. We are now in a position to make informed estimates of the sizes of the dimensionless groups appearing in (12)-(13). Both the Stokes number and the relative stiffness $\kappa$ depend on the forcing time scale $\mathcal{T}$. In particular, $\mathrm{St}\ll 1$ for $\mathcal{T}>1$, but the Stokes number may be large for faster movements. Similarly, $\kappa$ can acquire a large range of values. More usefully, the plate’s inertia $\varrho<10^{-3}$ for all physical values of $\mathcal{T}$ and hence may be safely neglected if required. Similarly, the reduced Reynolds number, $\epsilon^{2}\mathrm{Re}\approx 10^{-3}$ and fluid inertia may be neglected. Given the uncertainty in the modulus of the cupula itself, we note only that the inverse bendability, $b=B/(Ta^{2})$, is expected to be large.
The size of the dimensionless groups discussed above will inform our choices when neglecting terms in the governing equations (11)-(12). In particular, we will exploit the smallness of $\epsilon$ to neglect terms of order $\epsilon^{2}\mathrm{Re}$, and the smallness of $\varrho$ to neglect inertial terms in the plate equation (13). The last simplification is particularly useful as it will allow us to simplify spatio-temporal PDEs into two sets of ODEs.
We now turn to consider in more detail the behaviour of the model for the typical parameter values already discussed. Given the broad range of values that may be taken by the Stokes number, we begin by considering slow movement (i.e. large $\mathcal{T}$) in§ 3, and thus neglect terms of size $\mathrm{St}$. This will allow us to derive a reduced order equation (an ODE) for the deflection of the cupula that can be compared to numerical results in §4. However, in §5 we consider fast movements with finite Stokes numbers, leading to an integro-differential equation for the deflection of the cupula. Both the relative stiffness $\kappa$ and the inverse bendability $b$ are treated as independent parameters. In particular, recalling that $\kappa$ is a key parameter, capturing the relative timescales of the imposed motion and the cupular relaxation, a main objective of our analysis will be to investigate how variations in $\kappa$ give rise to different solution regimes.
3 Asymptotic solution for negligible Stokes number
Motivated by the small value of the Stokes number $\mathrm{St}=a^{2}/(\nu\mathcal{T})\sim 10^{-2}$ for natural movement timescales, $\mathcal{T}\sim 1$ sec, in this section we assume the Stokes number is negligibly small and solve the coupled system (12) asymptotically, expanding the solution in powers of the small aspect ratio $\epsilon$. To this end we introduce the following formal expansions
$\begin{split}u(r,\theta,s,t)&=u_{0}(r,\theta,s,t)+\epsilon u_{1}(r,\theta,s,t)+\epsilon^{2}u_{2}(r,\theta,s,t)+\cdots\\ v(r,\theta,s,t)&=v_{0}(r,\theta,s,t)+\epsilon v_{1}(r,\theta,s,t)+\epsilon^{2}v_{2}(r,\theta,s,t)+\cdots\\ w(r,\theta,s,t)&=w_{0}(r,\theta,s,t)+\epsilon w_{1}(r,\theta,s,t)+\epsilon^{2}w_{2}(r,\theta,s,t)+\cdots\\ p(r,\theta,s,t)&=p_{0}(r,\theta,s,t)+\epsilon p_{1}(r,\theta,s,t)+\epsilon^{2}p_{2}(r,\theta,s,t)+\cdots\\ \eta(r,\theta,t)&=\eta_{0}(r,\theta,t)+\epsilon\eta_{1}(r,\theta,t)+\epsilon^{2}\eta_{2}(r,\theta,t)+\cdots\end{split}$ (16)
Substitution of (16) into the dimensionless problem (12) yields a system of linear equations; we shall retain terms up to and including $\mathcal{O}(\epsilon)$ since they are required to explain the symmetry breaking phenomenon that was observed numerically (see fig. 3).
3.1 Expanded solution
Substitution of the formal expansion (16) into the governing equations (12), yields the following balance after discarding terms of order $\epsilon^{2}$ and higher:
  • $\mathcal{O}(1):$
    $f(t)$ $=-\kappa\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{0}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{0}}{\partial\theta^{2}},$ (17a)
    $\frac{\partial p_{0}}{\partial r}=\frac{\partial p_{0}}{\partial\theta}=0,$ (17b)
    $\frac{\partial u_{0}}{\partial r}+\frac{u_{0}}{r}+\frac{1}{r}\frac{\partial v_{0}}{\partial\theta}+\frac{\partial w_{0}}{\partial s}=0,$ (17c)
    $\nabla^{4}$ $\eta_{0}-b^{-1}\nabla^{2}\eta_{0}=\Delta p_{0},$ (17d)
  • $\mathcal{O}(\epsilon):$
    $f(t)\epsilon r\cos\theta$ $=-\kappa\frac{\partial p_{1}}{\partial s}+\kappa r\cos\theta\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{1}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{1}}{\partial\theta^{2}}+\cos\theta\frac{\partial w_{0}}{\partial r},$ (18a)
    $\kappa\frac{\partial p_{1}}{\partial r}=\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}\cos\theta,\quad\frac{\kappa}{r}\frac{\partial p_{1}}{\partial\theta}=-\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}\sin\theta,$ (18b)
    $\frac{\partial u_{1}}{\partial r}+\frac{u_{1}}{r}+\frac{1}{r}\frac{\partial v_{1}}{\partial\theta}+\frac{\partial w_{1}}{\partial s}=\cos\theta\left(r\frac{\partial w_{0}}{\partial s}-u_{0}\right),$ (18c)
    $\nabla^{4}\eta_{1}-b^{-1}\nabla^{2}\eta_{1}+\varrho r\cos\theta f(t)=\Delta p_{1}.$ (18d)
We identify the usual lubrication/boundary layer theory result that the leading order pressure is constant along a cross section (Craster & Matar, 2006). Furthermore, we note that the differential operators acting on the continuity and momentum equations are the same as in cylindrical coordinates (Batchelor, 1973).
Our starting point is the continuity equation, which can be integrated over a cross-section to deduce that the flux $Q=\int_{0}^{2\pi}\int_{0}^{a(s)}rw(r,\theta,s,t)\,\mathrm{d}r\,\mathrm{d}\theta$ is conserved in the $s$ direction, i.e. $\partial Q/\partial s=0$. This means that the flux is exclusively a function of time, a fact we will exploit to derive a reduced-order equation. Invoking symmetry we now seek an axisymmetric solution with $v_{0}=0$ and leading order terms independent of $\theta$.
Turning our attention to the $\mathcal{O}(\epsilon)$ problem, we note that $p_{1}(r,\theta,s,t)$ may be decomposed into a pressure gradient along the duct axis and an $s$-independent pressure variation due to centrifugal effects, so that
$p_{1}(r,\theta,s,t)=\bar{p}_{1}(s,t)+\frac{\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}}{\kappa}r\cos\theta,$ (19)
and as we only require the $s$ component of the pressure gradient in the computation of the axial velocity, we may safely ignore the $s$-independent component of $p_{1}$ and use $\bar{p}_{1}(s,t)$ in its place. Moreover, the first order pressure jump across the cupula $\Delta p_{1}=p_{1}(s=2\pi,t)-p_{1}(s=0,t)=\Delta\bar{p}_{1}$.
The solutions giving the first two orders of the axial velocity in terms of the pressure gradients may be determined directly: $w_{0}$ is found from (17a), under the assumption of axisymmetry, while $w_{1}$ may be found by decomposing it into axisymmetric and asymmetric parts and using standard methods. We find that
$w_{0}(r,s,t)=-\frac{1}{4}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)(a(s)^{2}-r^{2}),$ (20a)
$w_{1}(r,\theta,s,t)=-\frac{1}{4}\kappa\frac{\partial\bar{p}_{1}}{\partial s}(a(s)^{2}-r^{2})+\frac{1}{16}\left(f(t)-3\kappa\frac{\partial p_{0}}{\partial s}\right)r(r^{2}-a(s)^{2})\cos\theta.$ (20b)
Figure 4: Velocity profiles predicted by (20) as the torus aspect ratio, $\epsilon$, and leading-order term vary. Note how when $f(t)$ and $\kappa\Delta p/(2\pi)$ cancel each other, the asymmetric flow dominates. Moreover, the symmetry breaking becomes observable earlier for larger values of $\epsilon$, though we reiterate that our theory is formally valid only for $\epsilon\ll 1$.
We can get a first hint of the symmetry breaking mechanism observed in Figure 3 by considering when the asymmetric correction term $\epsilon w_{1}$ is of a similar size as the symmetric leading order solution $w_{0}$. Indeed, it is easy to verify that this will be the case when the pressure gradient approximately cancels out the forcing $f(t)$, such that the modulating coefficient $f(t)+\kappa\frac{\partial p_{0}}{\partial s}$ in (20a) is close to zero. This is visualized in Figure 4, where we plot the velocity $w=w_{0}+\epsilon w_{1}$ for several values of $\epsilon$ and $f(t)+\kappa\frac{\partial p_{0}}{\partial s}$, observing an asymmetric profile when $f(t)+\kappa\frac{\partial p_{0}}{\partial s}\ll 1$.
From the velocities (20), we may compute the flux, noting that the asymmetric components of $w_{1}(r,\theta,s,t)$ integrate to zero because of the $\cos\theta$ term:
$Q_{0}=-\frac{2\pi}{16}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)a(s)^{4},$ (21a)
$Q_{1}=-\frac{2\pi}{16}\frac{\partial\bar{p}_{1}}{\partial s}a(s)^{4}.$ (21b)
Since the flux $Q=Q_{0}+\epsilon Q_{1}+...$ is independent of $s$, we can now integrate the above equations along the axis of the duct, obtaining
$I_{4}Q_{0}=-\frac{\pi}{8}\left(2\pi f(t)+\kappa\Delta p_{0}\right),$ (22a)
$I_{4}Q_{1}=-\frac{\pi}{8}\Delta p_{1},$ (22b)
where we have defined $I_{4}=\int_{0}^{2\pi}a(s)^{-4}~{}\mathrm{d}s$, and used the fact that $\Delta p_{1}=\Delta\bar{p}_{1}$.
To connect the flow to the cupula displacement, we evaluate the flux using the velocity $w$ at the cupula’s location, where the fluid velocity must equal the velocity of the cupula: $w=\frac{\partial\eta}{\partial t}$. Therefore, we may write $Q_{0}=2\pi\int_{0}^{a_{0}}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r$ where $a_{0}=a(0)$. Moreover, substituting the pressure jump using the plate equations (17d) and (18d) yields the following equations for the displacement $\eta$:
$I_{4}\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}\mathrm{dr}=-\frac{1}{16}\left[2\pi f(t)+\kappa\left(\nabla^{4}\eta_{0}-b^{-1}\nabla^{2}\eta_{0}+\varrho f(t)\right)\right],$ (23a)
$I_{4}\int_{0}^{a(0)}r\frac{\partial\eta_{1}}{\partial t}\mathrm{dr}=-\frac{1}{16}\kappa\left(\nabla^{4}\eta_{1}-b^{-1}\nabla^{2}\eta_{1}+\varrho r\cos\theta f(t)\right).$ (23b)
(23) are equations that completely determine $\eta_{0}$ and $\eta_{1}$ when the Stokes number is negligible. Furthermore, as $\varrho\ll 1$, we may neglect it in (23). We may solve (23) by writing the deflection as $\eta=\Delta p(t)\bar{\eta}(r)$, where $\bar{\eta}$ satisfies $\nabla^{4}\bar{\eta}-b^{-1}\nabla^{2}\bar{\eta}=1$, from which we obtain a pair of ordinary differential equations for $\Delta p(t)$:
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\left(2\pi f(t)+\kappa\Delta p_{0}\right),$ (24a)
$\alpha(b)\frac{\mathrm{d}\Delta p_{1}}{\mathrm{d}t}=-\frac{1}{16I_{4}}\kappa\Delta p_{1}$ (24b)
where we have defined $\alpha(b)=\int_{0}^{a_{0}}r\bar{\eta}(r)\mathrm{d}r$. If we consider the cupula’s edges to be clamped, the solution for $\bar{\eta}$ is given by:
$\bar{\eta}(r)=\frac{b}{4}(a_{0}^{2}-r^{2})+\frac{a_{0}b^{3/2}}{2I_{1}(b^{-1/2}a_{0})}\left[I_{0}\left(b^{-1/2}r\right)-I_{0}\left(b^{-1/2}a_{0}\right)\right],$ (25)
with $I_{k}(\cdot)$ the modified Bessel function of the first kind and $k$th order; we then also have
$\alpha(b)=\int_{0}^{a_{0}}r\bar{\eta}(r)\mathrm{d}r=\frac{a_{0}^{2}b^{2}}{2}+\frac{a_{0}^{4}b}{16}-\frac{a_{0}^{3}b^{3/2}I_{0}\left(a_{0}b^{-1/2}\right)}{4\,I_{1}\left(a_{0}b^{-1/2}\right)}.$ (26)
We remark that when no pretension exists in the cupula $\alpha(b\rightarrow\infty)$ is well-defined and finite, and in particular $\alpha\rightarrow a_{0}^{6}/384$, so that $\alpha$ remains well defined in the plate limit. (24b) has an exponential solution, but the condition of homogeneous initial conditions implies that $\Delta p_{1}(0)=0$, from which we conclude $\Delta p_{1}(t)\equiv 0$. Therefore, there is no first order correction to the cupular pressure jump when $\mathrm{St}\ll 1$ and $\varrho\ll 1$, and the leading order pressure pressure jump
$\Delta p_{0}=-\frac{\pi}{8\alpha(b)I_{4}}\int_{0}^{t}f(\tau)e^{-\frac{\kappa}{16I_{4}\alpha(b)}(t-\tau)}\mathrm{d}\tau,$ (27)
is correct to $\mathcal{O}(\epsilon^{2})$.
3.1.1 Expressions for the velocities
Once the pressure jump is known from (24a), the leading order pressure gradient may be computed from
$\frac{\partial p_{0}}{\partial s}=\frac{\Delta p_{0}}{I_{4}a(s)^{4}}+\frac{1}{\kappa}\left[\frac{2\pi}{I_{4}a(s)^{4}}-1\right].$ (28)
Substitution into the axial velocity (20) then yields
$w_{0}$ $=\frac{\pi}{2I_{4}a(s)^{4}}\left[f(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right][r^{2}-a(s)^{2}],$ (29a)
$w_{1}$ $=\frac{r(r^{2}-a(s)^{2})\cos\theta}{16I_{4}a(s)^{4}}\left[-3\kappa\Delta p_{0}+(-6\pi+4I_{4}a(s)^{4})f(t)\right].$ (29b)
The radial and azimuthal velocities can be recovered from the continuity equation,
$u_{0}$ $=\frac{\pi}{2I_{4}a(s)^{5}}\left[f(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right]r[r^{2}-a(s)^{2}]\frac{\mathrm{d}a}{\mathrm{d}s},$ (30a)
$v_{0}$ $=0,$ (30b)
$u_{1}$ $=\frac{2f(t)\left[2I_{4}a(s)^{4}+\pi\right]+\kappa\Delta p_{0}}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\cos\theta\frac{\mathrm{d}a}{\mathrm{d}s},$ (30c)
$v_{1}$ $=-\frac{5\left[2f(t)\left(2I_{4}a(s)^{4}+\pi\right)+\kappa\Delta p_{0}\right]}{16I_{4}a(s)^{5}}r^{2}\left[r^{2}-a(s)^{2}\right]\sin\theta\frac{\mathrm{d}a}{\mathrm{d}s}.$ (30d)
Note that for a perfect i.e. uniform torus, $a^{\prime}(s)=0$ and there are no radial or azimuthal velocities.
3.2 Deformation regimes
For a fixed aspect ratio the key remaining dimensionless parameter is the relative stiffness $\kappa$ (which can be varied by changing the bending stiffness or timescale of forcing). We shall see that changing $\kappa$ makes the system sensitive either to the angular acceleration or the angular velocity. In this section we analyse how these limiting cases arise by solving (24a). Although (24a) can be solved analytically for any forcing through (27), qualitative information may be obtained by considering the large and small $\kappa$ limits.
When the cupula is soft ($\kappa\ll 1$) the solution to (24a) is approximately given by
$\Delta p\sim-\frac{\pi}{8\alpha(b)I_{4}}\int_{0}^{t}f(\tau)\mathrm{d}\tau=-\frac{\pi}{8I_{4}\alpha(b)}F(t)\quad\text{as}\quad\kappa\to 0.$ (31)
At the other extreme, for a stiff cupula, characterized by $\kappa\gg 1$, an approximate solution of (24a) is given by
$\Delta p\sim-\frac{2\pi}{\kappa}f(t)\quad\text{as}\quad\kappa\to\infty.$ (32)
In physical terms, the two results in (31) and (32) represent qualitatively distinct regimes for the response of the cupular displacement to the imposed rotation of the canal: for a soft cupula ($\kappa\ll 1$), the pressure difference across the cupula, and thus the cupula deformation, is proportional to the angular velocity of the imposed rotation, $F(t)$, while for a stiff cupula ($\kappa\gg 1$), the deformation instead follows the angular acceleration $f(t)=F^{\prime}(t)$. Since the cupular deformation is thought to be what is detected by the nerve cells in the cupula, this suggests that the cupula can detect either the angular velocity or the angular acceleration to which it is subject — depending on the value of $\kappa$.
In the latter case, it is easy to verify that the leading order radial and axial velocities suffer a cancellation, as the prefactor of the leading order velocity (20a) is $f(t)+\kappa\frac{\Delta p_{0}}{2\pi}=\mathcal{O}\left(\frac{1}{\kappa}\right)$, and the asymmetric order $\epsilon$ correction $w_{1}(r,\theta,s,t)$ dominates for any $\epsilon$ provided $\kappa$ is sufficiently large. This cancellation accounts for the behaviour observed in Figure 3: as the Young’s modulus is increased, making the cupula stiffer to the point that $\kappa\gg 1$, the flow ceases to be symmetric.
We emphasize that although symmetry breaking arises from the breakdown of the asymptotic ordering between the first and second terms in the series, this does not imply a loss of asymptotic ordering in the higher-order terms. The symmetry breaking results from a catastrophic cancellation in the leading-order term, while the first correction remains $\mathcal{O}(\epsilon)$, and the subsequent terms are expected to retain their anticipated scaling—indicating that the series remains well-behaved. We confirm this in the next section by numerically solving the full nonlinear problem (1).
To estimate the critical of $\kappa$ where the transition occurs, we may write $\Delta p_{0}=Ae^{2\pi it}+c.c.$ and $f(t)=Be^{2\pi it}+c.c.$, with $A,B\in\mathbb{C}$,and seek the range of $\kappa$ for which $\Delta p$ is in phase with $f$. Direct substitution into (24a) yields
$A=-\frac{\pi}{8I_{4}}\frac{B}{2\pi\alpha(b)i+\kappa/(16I_{4})}.$ (33)
Therefore, the transition change occurs at $\kappa_{c}=32I_{4}\alpha(b)$. For $\kappa<\kappa_{c}$ the phase difference between $A$ and $B$ is more than $\frac{\pi}{4}$, and it will be less than $\frac{\pi}{4}$ for $\kappa>\kappa_{c}$. When the phase difference is small the response is roughly in phase with the forcing $f(t)$, and vice-versa.
3.2.1 Sample head rotation
To visualize the different regimes we have identified above, we solve the equation for the pressure jump (24a) with a specific choice for the forcing $f(t)$. While many forms of $f(t)$ could be considered, we are motivated by a clinical head manoeuvre that may be thought of as modelling a slow rotation of the head from right to left. Although other clinical models are described by high-order polynomials (Boselli et al., 2009, 2013), we choose a particular form that facilitates an analytical solution of the equations, namely:
$f(t)=\frac{\mathrm{d}F}{\mathrm{d}t}=\begin{cases}\sin 2\pi t&t\in(0,1)\\ 0&t>1.\end{cases}$ (34)
Solving for the pressure gradient through (27) yields
$-\Delta p=\beta\int_{0}^{t}f(\tau)e^{-\gamma\kappa(t-\tau)}\mathrm{d}\tau=\begin{cases}\beta\frac{2\pi e^{-\gamma\kappa t}+\gamma\kappa\sin 2\pi t-2\pi\cos 2\pi t}{\gamma^{2}\kappa^{2}+4\pi^{2}}&t<1,\\ \frac{2\pi\beta}{4\pi^{2}+\gamma^{2}\kappa^{2}}e^{-\gamma\kappa t}(1-e^{-\gamma\kappa})&t>1.\end{cases}$ (35)
Here, $\beta=\pi/[8I_{4}\alpha(b)]$ and $\gamma=\kappa/[16\alpha(b)I_{4}]$. The cupular deformation, which in this regime is proportional to the pressure jump, is plotted in the left panel of Figure 5 for different values of $\kappa$, from which we can clearly see the transition from $\Delta p$ tracking the angular velocity $F(t)$ for small $\kappa$ to tracking the angular acceleration $f(t)$ for large $\kappa$, as expected from the preceding analysis. There is an interesting transition region for $\kappa\sim 1$, where we can see an “overshoot” region at the end of the manoeuvre that has not decayed. This is not the case for either of the limiting regions, where the pressure jump (and cupular displacement) is identically zero after the completion of the head turn.
Figure 5: The influence of dimensionless stiffness $\kappa$ on the cupular deformation. (a) As $\kappa$ is increased, the deformation (normalized by the maximum) transitions from following the angular velocity to following the angular acceleration. (b) This transition with $\kappa$ may be shown by plotting the correlation, $R$, between the deformation and the angular velocity $\Omega(t)$ (solid curve) or the angular acceleration $\dot{\Omega}(t)$ (dashed curve). A transition between the two regimes occurs around $\kappa\approx 100$. In both plots, colour is used to show the value of $\kappa$.
In Figure 5(b) we compare how similar the response is to either the angular velocity or the angular acceleration by computing the correlation between the respective functions; for two functions $f(t)$ and $g(t)$, this correlation is defined as
$R(f,g)=\frac{\int_{0}^{T}f(t)g(t)~{}\mathrm{d}t}{\left(\int_{0}^{T}f(t)^{2}~{}\mathrm{d}t\cdot\int_{0}^{T}g(t)^{2}~{}\mathrm{d}t\right)^{1/2}}.$ (36)
As expected from our asymptotic analysis, $\Delta p$ correlates with the angular velocity for small to moderate values of $\kappa$, and the angular acceleration for large values of $\kappa$. For the parameters used in Figure 5, we compute $\kappa_{c}=32\pi\approx 100$, in agreement with the transition point observed in the plot.
4 Numerical simulations
To test the validity of our asymptotic approach, we return to the numerical simulations in COMSOL as presented in §2.2 but now imposing within the numerical scheme the forcing given by (34), and varying the Young’s modulus of the solid material to change $\kappa$. We perform two direct comparisons, appearing in Figures 6 and 7. Figure 6 plots the cupular pressure jump $\Delta p$ as a function of time, while Figure 7 plots the axial velocity profiles across the cross section, sampled at several time points for different values of $\kappa$ in a region of the canal far from the cupula. In both figures, we observe excellent agreement between theory and numerics. In particular, the breaking of symmetry in the flow profile for large $\kappa$ is easily observed in Figure 7, and the theoretical profile captures the trend and profile shape very well.
[…]
We remark that although the symmetry breaking might suggest a breakdown of the asymptotic order, with the second term dominating the first in the series, subsequent terms are well behaved and the series is not divergent. This may be inferred from the agreement between the model solution and the numerical solution to the full nonlinear problem, even for large $\kappa$ values where the symmetry is broken. Moreover, as the symmetry breaking occurs because the leading order term shrinks and the correction retains its size (rather than growing), we expect higher order terms to retain their sizes too, preserving the asymptotic order of the solution.
We have thus far considered slow movements, so that $\mathcal{T}>1\mathrm{~{}s}$ and $\mathrm{St}\ll 1$. However, for faster movements, typically when $\mathcal{T}<1\mathrm{~{}s}$, the Stokes number is no longer negligible and inertial terms must be retained in the analysis (see §2.4). However, the Stokes number may also become non-negligible for other reasons — for example, some authors report slightly thicker semicircular canals ($a$ slightly larger), so that the Stokes number is considerably larger than expected due to its quadratic dependence on velocity. To this end, in the next section we consider flows with a finite Stokes number.
Figure 6: Comparison of the pressure difference across the cupula as predicted by COMSOL simulations (markers) and the theoretical prediction from (35) (Solid line). As expected from the results in Section 3.2, depending on the value of $\kappa$ the deformation tracks the angular velocity or angular deformation of the forcing (given by (34)). The parameter values used are given in the main text, which correspond to $\mathrm{St}=0.0256$.
Figure 7: The numerically obtained velocity profiles (blue markers) and theoretical predictions (black solid curves) from (20a) and (20b). As $\kappa$ increases, the velocity profile ceases to be symmetric around $\kappa\approx 10^{3}$. The imposed rotation is given by (34), sampled at 7 different times. Parameter values are the same as in fig. 6.
5 Effect of fluid inertia
While considering the inertialess limit of $\mathrm{St}\ll 1$ facilitated analytical progress, there are several circumstances in which inertia may become important, e.g. faster movements or larger canals. Therefore, we consider the effect of fluid inertia by retaining the $\mathcal{O}(\mathrm{St})$ terms in the governing equations (17a). Proceeding as in the previous section and seeking a solution in powers of the small aspect ratio $\epsilon$, we seek a solution for the leading order axial velocity of the form
$w_{0}(r,s,t)=\sum_{n=1}^{\infty}c_{n}(t,s)\phi_{n}(r,s),\quad\phi_{n}(r,s)=J_{0}\left(\frac{\lambda_{n}}{a(s)}r\right).$ (37)
Here, $\phi_{n}$ are the eigenfunctions for the Laplacian in a cylinder of local radius $a(s)$, subject to Dirichlet boundary conditions (Batchelor, 1973); the $\lambda_{n}$ are the zeros of the Bessel function of the first kind $J_{0}(z)$, thereby ensuring that the axial velocity satisfies the no-slip condition at $r=a(s)$ in the rotating frame. Thus, substituting (37) into the momentum equation, and using the orthogonality properties of Bessel functions we find
$\operatorname{St}\frac{\partial c_{n}}{\partial t}=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)-\frac{\lambda_{n}^{2}}{a(s)^{2}}c_{n},$ (38a)
$c_{n}(t,s)=-\frac{2}{\lambda_{n}J_{1}(\lambda_{n})}\int_{0}^{t}\left(f(\tau)+\kappa\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~{}\mathrm{d}\tau,$ (38b)
where $\mathcal{K}_{n}(x,s;\mathrm{St})=\mathrm{St}^{-1}e^{-\lambda_{n}^{2}x/[a(s)^{2}\mathrm{St}]}$. The flux may now be computed as
$\begin{split}Q_{0}&=2\pi\int_{0}^{a(s)}rw_{0}(r,s,t)~{}\mathrm{d}r=2\pi a(s)^{2}\int_{0}^{1}\rho w_{0}(a(s)\rho,s,t)~{}\mathrm{d}\rho\\ &=-4\pi a(s)^{2}\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)\mathcal{K}_{n}(t-\tau,s;\mathrm{St})~{}\mathrm{d}\tau\right].\end{split}$ (39)
From the continuity equation, the flux $Q_{0}$ is independent of $s$. Therefore, we can evaluate it at the location of the cupula, where the fluid velocity is known to be equal to the cupular velocity $\frac{\partial\eta_{0}}{\partial t}$; this gives $Q_{0}=2\pi\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}~{}\mathrm{d}r$ and allows us to write a reduced system of equations for $\frac{\partial p_{0}}{\partial s}(s,t)$ and $\eta_{0}(r,t)$, namely
$\int_{0}^{a(0)}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r=-2a^{2}\frac{1}{\operatorname{St}}\sum_{n=1}^{\infty}\left[\frac{e^{-\lambda_{n}^{2}t/(a^{2}\operatorname{St})}}{\lambda_{n}^{2}}\int_{0}^{t}\left(f(t)+\kappa\frac{\partial p_{0}}{\partial s}\right)e^{\lambda_{n}^{2}\tau/(a^{2}\operatorname{St})}~{}\mathrm{d}\tau\right],$ (40a)
$\nabla^{4}\eta_{0}-\nabla^{2}\eta_{0}+\varrho\left(\mathrm{St}\frac{\partial^{2}\eta_{0}}{\partial t^{2}}+f(t)\right)=\Delta p_{0}=\int_{0}^{2\pi}\frac{\partial p_{0}}{\partial s}~{}\mathrm{d}s.$ (40b)
For an arbitrary inner radius $a(s)$, (40) can be solved using the Laplace transform, as shown in Appendix C. Before tackling the general case, we focus on the simple case where the tube radius is uniform, $a(s)\equiv 1$.
5.1 A simple example
For the special case when $a(s)\equiv 1$, i.e. the tube radius is constant, the pressure gradient can be assumed to be independent of $s$, that is $\frac{\partial p_{0}}{\partial s}=\frac{\Delta p_{0}}{2\pi}$, and (40a) simplifies to
$\int_{0}^{1}\frac{\partial\eta_{0}}{\partial t}r\mathrm{d}r=-2\sum_{n=1}^{\infty}\left[\lambda_{n}^{-2}\int_{0}^{t}\left(f(\tau)+\kappa\frac{\Delta p_{0}}{2\pi}\right)\mathcal{K}_{n}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau\right].$ (41)
To transform this equation into a more manageable form, we define a complete kernel $\mathcal{K}(x;\mathrm{St})=\mathrm{St}^{-1}\sum_{n=1}^{\infty}\lambda_{n}^{-2}e^{-\lambda_{n}^{2}x/\mathrm{St}}$. Finally, we may directly substitute $\Delta p_{0}$ from (40b) into (39) to obtain a single equation for the cupular deflection $\eta_{0}(r,t)$:
$\int_{0}^{1}\frac{\partial\eta_{0}}{\partial t}r~{}\mathrm{d}r=-2\bar{f}(t)-\frac{\kappa}{\pi}\int_{0}^{t}\left[\nabla^{4}\eta_{0}-b^{-1}\nabla^{2}\eta_{0}+\varrho\left(\mathrm{St}\frac{\partial^{2}\eta_{0}}{\partial t^{2}}+f(\tau)\right)\right]\mathcal{K}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau,$ (42)
where we have introduced $\bar{f}(t)=\int_{0}^{t}f(\tau)\mathcal{K}(t-\tau;\mathrm{St})~{}\mathrm{d}\tau$.
Equation (42) is an integro-PDE for $\eta$ with an exponential-like kernel. Seeking a separable solution $\eta_{0}=\bar{\eta}(r)\Delta p_{0}$ when $\varrho\ll 1$ will lead to an integro-differential equation for $\Delta p_{0}(t)$
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2\bar{f}-\frac{\kappa}{\pi}\int_{0}^{t}\Delta p_{0}(\tau)\mathcal{K}(t-\tau;\mathrm{St})\mathrm{d}\tau.$ (43)
Here, $\bar{\eta}$ and $\alpha(b)$ are the same as in the $\mathrm{St}\ll 1$ limit, given in (25). Equation (43) may be efficiently solved numerically by truncating the infinite series in the kernel and transforming the integral equation into a system of ODEs. This is a standard calculation, with details given in Appendix B.
5.2 Fluid inertia can make the cupula underdamped
To understand the effect of inertia in the cupular response, we first consider the case of a cupula that is initially stretched by a pressure jump $\Delta p_{0}(t=0)=1$ in a frame rotating at constant speed, so that $f(t)=0$.
The numerical solution to (43) for a range of Stokes numbers is given in Figure 8(a). Notice that for sufficiently large $\mathrm{St}$, $\Delta p(t)$ exhibits decaying oscillatory behaviour, meaning that the cupula is underdamped; this is in contrast to smaller values of $\mathrm{St}$, in which the cupula dynamics show an exponential decay whose rate of decay increases with $\mathrm{St}$.
Figure 8: Left panel: Solution to (43), when $f(t)=0$ and $\Delta p_{0}(t=0)=1$ for different values of the Stokes number, showing underdamped dynamics for large enough $\mathrm{St}$. Central panel: Bifurcation diagram, showing the evolution of $\mathrm{Re}(\bar{\omega})$ (blue) and $\mathrm{Im}(\bar{\omega})$ (red). Markers represent the numerically obtained solution from (44) and dashed lines the analytical approximation (45). Right panel: bifurcation diagram for $f(t)\sim e^{2\pi it}$, showing how the transition between the two regimes $\Delta p_{0}\sim F(t)$ and $\Delta p_{0}\sim f(t)$ depends on both $\kappa$ and $\mathrm{St}$. Colour represents the complex angle of $\chi$, with blue $\arg(\chi)=\pi/2$ and red representing $\arg(\chi)=0$.
To understand this change in behaviour as $\mathrm{St}$ is increased, we seek an exponential ansatz to solve (43), with $\Delta p_{0}\sim e^{-\omega t}$ in the simple case when $f(t)=0$. Direct substitution yields the following condition:
$\frac{\alpha(b)\pi}{\kappa\mathrm{St}}\bar{\omega}=\sum_{n=0}^{\infty}\frac{1}{\lambda_{n}^{2}}\frac{1}{\lambda_{n}^{2}-\bar{\omega}},$ (44)
where the rescaled growth rate is $\bar{\omega}=\mathrm{St}\,\omega$. Equation (44) provides an equation for $\bar{\omega}$ depending on the parameter $q=\alpha\pi/(\kappa\mathrm{St})$. Analytical progress can be made by truncating the sum at $n=0$, i.e. considering only the first term. This leads to
$\bar{\omega}=\frac{\lambda_{0}^{2}}{2}\left[1\pm\left(1-\frac{4}{q\lambda_{0}^{6}}\right)^{1/2}\right].$ (45)
From this, we can identify the critical value for the parameter $q$ at which the transition to underdamped dynamics occurs: $q_{c}=4/\lambda_{0}^{6}$.
A natural question to ask now is if fluid inertia can alter the development of the two flow regimes outlined previously in Section 3.2? Proceeding as before, we assume a forcing $f(t)=Be^{2\pi it}$, and try an ansatz $\Delta p_{0}=Ae^{2\pi it}$. Substitution into the integral equation (43) and neglecting contributions from the initial conditions yields
$2\pi i\alpha A=\sum_{n=0}^{\infty}\frac{2B+A\pi\kappa}{\lambda_{n}^{2}(2\pi i\mathrm{St}+\lambda_{n}^{2})}$ (46)
It is convenient to define the function $\mathcal{G}:\mathbb{R}\rightarrow\mathbb{C}$, given by $\mathcal{G}(\mathrm{St})=\sum_{n=0}^{\infty}\lambda_{n}^{-2}/(2\pi i\mathrm{St}+\lambda_{n}^{2})$. We find that the response (characterized by $A$) is related to the forcing (characterized by $B$) through
$A=\frac{-2B\mathcal{G}(\mathrm{St})}{2\pi i\alpha+\pi\kappa\mathcal{G}(\mathrm{St})}.$ (47)
Therefore, we see that the angle of the complex quantity
$\chi=\frac{\mathcal{G}(\mathrm{St})}{2\pi i\alpha+\pi\kappa\mathcal{G}(\mathrm{St})}$ (48)
will determine if the deformation follows the angular velocity (if the angle is close to $\pi/2$) or the angular acceleration (when the angle is close to 00 or multiples of $\pi$). To achieve analytical progress we truncate the sum in $\mathcal{G}$, keeping only the first term, and we find
$\mathrm{Re}(\chi)=\pi\kappa-4\pi^{2}\mathrm{St}\alpha\lambda_{0}^{2},\quad\mathrm{Im}(\chi)=2\pi\alpha\lambda_{0}^{4}.$ (49)
Therefore, the curve in the parameter space $(\kappa,\mathrm{St})$ separating the two regimes is given implicitly by
$\left|\frac{\kappa}{2\lambda_{0}^{4}\alpha}-\frac{2\pi\mathrm{St}}{\lambda_{0}^{2}}\right|=1.$ (50)
In Figure 8(c), we plot the argument of $\chi$ as a function of $\kappa$ and $\mathrm{St}$, indicating as well the approximate bifurcation curves given by (50). This diagram indicates where in the $\kappa$-$\mathrm{St}$ phase space the cupula deflection follows the angular velocity versus the angular acceleration. For small values of the Stokes number $\mathrm{St}$ we recover the previous picture, where $\kappa\ll 1$ indicates the deformation follows the angular velocity and $\kappa\gg 1$ indicates the response is guided by the angular acceleration. But we also find a nonlinear dependence on $\mathrm{St}$ for non-small Stokes numbers. For given $\kappa$ less than about 100, a transition to angular acceleration tracking occurs for $\mathrm{St}$ greater than about 1, meaning that the cupula system may be tuned to follow angular acceleration even for small cupula stiffness if the Stokes number is high enough. This transition point increases for larger $\kappa$, as indicated by the blue wedge region in Figure 8(c), meaning that an orders of magnitude higher $\kappa$ is possible with cupular deflection still tracking angular velocity, if the Stokes number is accordingly increased in a very particular way.
We may interpret the effect of high fluid inertia by considering the response of the cupula as the forcing frequency is increased. For small forcing frequencies the deformation will follow the angular acceleration, and as the forcing frequency is increased the cupula will start deforming in phase with the angular velocity, as expected from § 3.2; this is well known in the vestibular literature (Benson, 1990). However, our results from this section suggest that when the forcing frequency is further increased (in humans, to about 100 Hz), the cupular deformation will again be in phase with the angular acceleration.
5.3 Non-uniform channel widths
As noted in the introduction, an advantage of our theoretical formulation is that it is also compatible with a non-uniform and arbitrary channel width, described by the function $a(s)$, so long as the small aspect ratio between channel width and length is maintained. This case is more delicate than the one we saw in the last subsection, as the pressure gradient is no longer constant but depends on $a(s)$, and must be integrated along the channel to obtain the pressure jump across the cupula. In Appendix C a method based on the Laplace transform and the convolution theorem is developed, through which we obtain an approach to solving the problem for both non-uniform channel widths and $\mathrm{St}>0$. The main result is that we obtain the same form of equation (42) for the deflection $\eta_{0}(r,t)$, with the Kernel $\mathcal{K}(x;\mathrm{St})$ given by the (temporal) inverse Laplace transform of
$\tilde{\mathcal{K}}(\sigma;\mathrm{St})=\left(\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}\sum_{n=0}^{\infty}\lambda_{n}^{-2}\left[a(s)^{2}\mathrm{St}\sigma+\lambda_{n}^{2}\right]^{-1}}\right)^{-1}.$ (51)
For a given canal profile $a(s)$, the Kernel may be numerically obtained by fixing a discretization of $\sigma$ into a finite number of points. For each point, the integral can be populated for $a(s)$ and computed using standard quadrature methods. This will yield $\tilde{\mathcal{K}}(\sigma;\mathrm{St})$ for a finite number of $\sigma$. The equation for the pressure jump (43), can then be either solved in Laplace space, inverting the transformed solution using an efficient algorithm (Kuhlman, 2013), or in real space using a trapezoidal method.
We have developed a general framework that allows us to solve for the cupular displacement and pressure jump for complicated canal geometries allowing also for the possibility of fluid inertia. As an example of the scenarios in which this approach might be useful, we now reconsider some of the numerical results presented in figure 4.
5.4 Fluid inertia explains discrepancies between numerics and model
Figure 4 generally shows excellent agreement between the $\mathrm{St}=0$ model presented in § 3 and our COMSOL numerics, especially at the scale of the largest velocities, which were used for comparison in figure 4. However, if we zoom in to situations where the velocity is small, such as when $t=0.5$, then we might expect to observe differences caused by small errors in the phase of the motion. Figure 9 shows just such an effect: the agreement between the predictions of the $\mathrm{St}=0$ asymptotics (dashed black curves) and COMSOL simulations (blue markers) is no longer satisfactory for small and moderate values of the stiffness $\kappa\lesssim 1$: while the absolute error is small, the relative error is very large. This is because even for small Stokes numbers, the exact time at which the velocity zero is different from that predicted by the $\mathrm{St}=0$ asymptotics of § 3, which essentially assume that the motion is quasi-steady.
Figure 9 also shows the predictions of the leading-order in $\mathrm{St}$ asymptotic results presented in this section. As might be expected, we see much better agreement between the prediction accounting for $\mathrm{St}>0$ via (38b) and (43) (shown by solid red curves in figure 9) and the results of COMSOL simulations (points) than with the earlier result, which neglected the effects of fluid inertia entirely. We emphasize that this finite inertia case requires the numerical calculation of the integrals in (38b) and (43).
Figure 9: Velocity profiles for $t=0.5$, including the finite fluid inertia correction. Dashed black line: $\mathrm{St}=0$, blue markers: numerics, solid red line: $\mathrm{St}>0$ model.
We now consider if domain irregularities can give rise to interesting flow phenomena, in particular unexpected symmetry breaking and vortical flows when $\kappa\ll 1$, which have been reported previously (Boselli et al., 2013).
6 Analytical description of vortical flow
As a final point of analysis, we turn our attention to the possibility of vortical flow: several authors have reported the existence of vortical structures in computational studies of flow in the utricle, the enlarged portion of the semicircular canal (see Boselli et al., 2009; Goyens et al., 2019, for example). Vortices appear to occur even in the analogue of our limit $\kappa\ll 1$, when the flow in the slender portion of the canals is largely symmetric. In this section we use our asymptotic analysis to give an analytical description of such vortical flow structures and determine the geometrical conditions required for their emergence.
Mathematically, the reason why we might expect enlarged regions of the canal to experience symmetry breaking may be seen from the form of the leading order and $O(\epsilon)$ axial velocity terms in (29). In particular, in regions where $a(s)>1$, the magnitude of the leading order velocity $w_{0}$ is proportional to $1/[I_{4}\,a(s)^{4}]$, while the correction is of order $\epsilon$ the $f(t)$ term in (29b) remains $O(\epsilon)$ even as $a(s)^{4}$ becomes large. (We retain the $I_{4}$ term here, since it is also dependent on $a(s)$.) Hence, we might expect noticeable asymmetry in the flow to develop in regions where $1/[I_{4}a(s)^{4}]$ is comparable to $\epsilon$. Note that this can be achieved in this way without requiring that we are in the stiff cupula regime $\kappa\gg 1$ that led to the symmetry-breaking discussed in Section 3.2.
To demonstrate this possibility, we consider the predictions of our asymptotic theory for a canal with a localized bulge by taking the radius $a(s)$ to be the sum of a constant and a Gaussian:
$a(s)=1+(a_{m}-1)e^{-\gamma(s-\pi)^{2}}.$ (52)
Here, the cupula and utricle are located at $s=\pi$, $\gamma$ is a parameter controlling the width of the enlargement and $a_{m}$ is the maximum inner radius of the tube. This choice is motivated by the qualitative agreement with the imaging from Daocai et al. (2014). In Figure 10 we plot the velocity distribution for several channel geometries, in particular a top view of the mid plane of the flow around the canal, with colour indicating the magnitude of the axial velocity. The size of the enlarged region is increased from left to right (with $\gamma=1$ and $a_{m}=1,2,3,4$), and the appearance of the vortical flow is clear. The forcing is given by (34) and we visualize the solution at time $t=0.25$. Here we have used a small value of $\kappa=0.1$ so that the flow in the slender regions is predicted to remain largely symmetric, a feature that we will verify below.
Figure 10: Analytical reconstruction of vortical flow in the utricle as the maximum channel radius, $a_{m}$, increases. The channel has a largely uniform radius but is wider in the vicinity of the utricle — see (52) for the detailed profile of the tube. Here, we observe how as the size of the utricle is augmented the vortex develops. The forcing is given by (34) and we show the solution at time $t=0.25$. The parameters used are $\epsilon=0.05$, and $\kappa=0.1$. Furthermore we use the solution from § 3 that assumes the fluid inertia is vanishingly small, $\mathrm{St}=0$.
6.1 Conditions required for the formation of the utricular vortex
As noted above, symmetry breaking in the utricle occurs when the first order correction $\epsilon w_{1}$ is comparable with the leading order velocity $w_{0}$. Within a cross section, the maximum value of $w_{0}$ is attained at $r=0$, and is given by
$|w_{0}^{*}(s,t)|=\frac{\pi}{2I_{4}a(s)^{2}}\left|(t)+\frac{\kappa\Delta p_{0}}{2\pi}\right|\sim\frac{\pi}{2I_{4}a(s)^{2}}|(f(t)|),$ (53)
where the last approximation follows when $\kappa\ll 1$. The maximum value of $w_{1}$ is attained at $r=\frac{a(s)}{\sqrt{3}}$ and $\theta=0,\pi$, and is given by
$|w_{1}^{*}|=\frac{1}{24\sqrt{3}I_{4}a(s)}\left|-3\kappa\Delta p_{0}+\left(-6\pi+4I_{4}a(s)^{4}\right)f(t)\right|.$ (54)
Focusing on the utricle, where $a(s)$ is largest, $|w_{1}^{*}|$ is dominated by
$|w_{1}^{*}|\sim\frac{a_{m}^{3}}{6\sqrt{3}}\left|f(t)\right|.$ (55)
Comparing $w_{0}$ and the next term in the expansion, $\epsilon w_{1}$, we conclude that noticeably asymmetrical flow in the utricle first emerges when
$\frac{\pi}{2I_{4}a_{m}^{2}}\sim\epsilon\frac{a_{m}^{3}}{6\sqrt{3}},\Longleftrightarrow\frac{3\sqrt{3}\pi}{\epsilon}\sim a_{m}^{5}\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}}.$ (56)
The transition can be seen qualitatively in the left panel of Figure 10, where the flow in the utricle (left half of the torus) transitions from symmetric to asymmetric as the bulge is increased. In Figure 11 this transition is analysed quantitatively. In the left panel, the flow profile in the utricle is plotted for varying $a_{m}$. Observe that for large $a_{m}$, the flow utricle flow is asymmetric while the flow in the slender part of the canal remains symmetric. To quantify the transition to noticeably asymmetrical flow, we compute the correlation between the velocity profile $w=w_{0}+\epsilon w_{1}$ and the symmetric ($w_{0}$) and asymmetric ($w_{1}$) solutions. The correlations may be written as
$R(w,w_{0})=\int_{0}^{a}w(r)w_{0}(r)\mathrm{d}r=\sqrt{\frac{R_{0}}{R_{0}+\epsilon^{2}R_{1}}},$ (57a)
$R(w,w_{1})=\int_{0}^{a}w(r)w_{1}(r)\mathrm{d}r=\sqrt{\frac{R_{1}}{R_{0}+\epsilon^{2}R_{1}}},$ (57b)
$R_{0}=\int_{-a}^{a}w_{0}^{2}\mathrm{d}r=\frac{1}{15a_{m}^{3}I_{4}^{2}}\left(\kappa\Delta p_{0}+2\pi f(t)\right)^{2},$ (57c)
$R_{1}=\int_{-a}^{a}w_{1}^{2}\mathrm{d}r=\frac{1}{1680I_{4}^{2}a_{m}}\left(3\pi\Delta p_{0}+(6\pi-4I_{4}a_{m}^{4})f(t)\right).$ (57d)
Here, we have used the fact that $\int_{-a}^{a}w_{0}w_{1}\mathrm{d}r=0$, which is trivial since $w_{0}w_{1}$ is an odd function of $r$. The correlation $R(w,w_{0})$ will be close to 1 when the flow is largely symmetrical, and closer to zero when the flow is noticeably asymmetrical. The opposite is true of the correlation $R(w,w_{1})$. In the the right panel of fig. 11 we plot the symmetrical correlation $R(w,w_{0})$ (solid line) and the asymmetrical correlation $R(w,w_{1})$ (dashed line) as predicted from our analytical model.
Based on the definition of $\xi$ we expect a transition when $\xi\sim 1$. Fig. 11 confirms this expectation, showing that a transition indeed occurs at this point when the geometry $a(s)$ satisfies (52). We note that for times when $f(t)=0$, our analysis does not apply, and the flow remains symmetric, even for large utricles. This follows from (55), as it is clear that $w_{1}=0$ when $f(t)=0$.
Our analysis of the onset of vortical flow is further supported by numerical simulations, represented by triangles and stars in the righthand panel of fig. 11. These simulations were conducted using COMSOL, modelling the utricle as an ellipsoidal expansion of the toroidal geometry. We find good agreement between the analytical predictions and COMSOL simulations regarding the emergence of vortical flow. (We attribute the small discrepancy between the COMSOL and analytical model to the fact that the geometry is not identical in both cases.) Furthermore, the consistency between the numerical simulations and analytical solutions suggests that, although the first-order correction may be larger than the leading-order term, the subsequent terms in the series remain well behaved, ensuring that the asymptotic ordering is preserved.
Figure 11: Left panel: analytical reconstruction of flow profiles in the wide region of the channel (representing the utricle), $w(r,\theta,s=\pi,t)$ for different values of the maximum enlargement $a_{m}$. The inset shows the flow profiles in the thin region of the flow $w(r,\theta,s=0,t)$; these remain symmetric, confirming the symmetry breaking mechanism is not the same as the global symmetry-breaking mechanism discussed in § 3.2. Right panel: correlation (as defined in (57)) between the axial velocity in the utricle $w(r)=w_{0}(r)+\epsilon w_{1}(r,\theta)$ and the symmetric (solid) and asymmetric flow profile (dashed). Curves show the results of the analytical computation and triangles and stars show the correlations computed from the COMSOL solution. We find that the transition occurs when $\xi=a_{m}^{5}I_{4}\epsilon/(3\sqrt{3}\pi)\sim 1$, as predicted by our analysis.
7 Discussion and conclusion
In this study, we developed a mathematical framework to model fluid flow in the semicircular canals of the vestibular system, focussing in particular on the interaction between the fluid motion and cupular deformation. Through a systematic analytical and numerical investigation, we identified distinct physical regimes and key mechanisms that govern the fluid-structure response to an imposed rotation. Our results not only advance the understanding of flow dynamics in these biologically relevant systems but also provide a simple framework with the potential for analysing vestibular function and dysfunction in response to head movements.
Our analytical approach consisted in solving the Navier-Stokes equations via an asymptotic series in the small aspect ratio of the semicircular canals. Through asymptotic analysis, and by connecting the fluid flow at the cupula to the cupular deformation, described via a plate equation, we reduced the vestibular dynamics to an ODE system for the cupular deflection, whose behaviour could easily be characterized. In this way, we established two primary regimes of flow-cupula interaction, depending on the value of the relative stiffness parameter $\kappa$:
  • Soft cupula regime, $\kappa\ll 1$: When the cupular stiffness is relatively low, the deformation of the cupula closely follows the angular velocity of the head. In this regime, the flow in the canal exhibits symmetry about the centreline.
  • Stiff cupula regime, $\kappa\gg 1$: As the stiffness of the cupula increases, the deformation transitions to follow the angular acceleration of the head. In this regime, the symmetry of the fluid flow about the centreline is broken, creating distinct zones of differential flow. This transition highlights the importance of structural properties of the cupula in shaping the dynamic response of the vestibular system.
To verify the analytical findings, we conducted numerical simulations of the reduced equations using COMSOL Multiphysics. The numerical results showed excellent agreement with the asymptotic predictions, confirming the validity of the analytical approximations across a wide parameter space. Importantly, the numerical approach enabled us to explore the influence of complex fluid–solid interaction boundary conditions and the nonlinear, advection, term in the Navier-Stokes equations that are otherwiseintractable analytically.
When the inertial terms were incorporated into the governing equations via inclusion of a finite Stokes number, $\mathrm{St}$, we observed important modifications to the system’s behaviour. For small Stokes numbers, and in studying the relaxation of the cupula to an initial deformation, the system exhibited overdamped dynamics. This behaviour is also predicted when inertia is neglected, and is consistent with the low-Reynolds-number assumption inherent to the vestibular fluid dynamics. However, for sufficiently large Stokes numbers, the system exhibited underdamped oscillations, with the cupular deformation following angular acceleration even in the soft cupula regime. This transition highlights the interplay between inertial and viscous forces in shaping the dynamic response of the system. Physiologically, this finding suggests that under certain conditions, such as during rapid head movements, the vestibular system may exhibit enhanced sensitivity to acceleration due to inertial effects.
The assumption of an idealized toroidal geometry allowed for significant analytical simplifications but is also a significant simplification of the true anatomy of the semicircular canals, which exhibit variations in cross-sectional shape. To address this, we extended our analysis to more realistic geometries, focusing on domains with a single enlarged region that deviates from the perfect torus. In these regions, we found thatsignificant deviation in radius gives another mechanism for the breaking of flow symmetry (in addition to the rigidity–induced effect already discussed). These results provide new insights into the functional implications of anatomical variability in the semicircular canals. For example, variations in canal geometry across species or due to developmental differences may influence the sensitivity and response characteristics of the vestibular system.
In each of the scenarios considered, our analytical approach enabled us to derive explicit expressions for the transition point between physical regimes, that is we obtained formulas for the critical values of the relevant system parameters at which the transition between different regimes occurs. These formulas lend insight into the fine balance between different components of the system, and enable to speculate on how the vestibular system may have been fine tuned by evolution in different organisms, and/or key considerations in engineering an artificial vestibular system. We turn to such considerations next.
7.1 Implications and applications
The findings of this study have several implications for both biology and engineering. In the context of vestibular physiology, our results contribute to a deeper understanding of how the semicircular canals transduce head motion into neural signals. The distinction between velocity-sensitive and acceleration-sensitive regimes suggests that while the mechanical properties of the cupula, combined with canal geometry, enable the system to function under a wide range of motion frequencies (Bronstein et al., 2013) it emphasizes that what is sensed differs markedly across this parameter space. This flexibility of sensing may be useful for maintaining balance and spatial orientation across diverse locomotor activities (Golding & Gresty, 2005, 2015).
Here, it is worth considering the distinction between the soft cupula ($\kappa\ll 1$) and stiff cupula ($\kappa\gg 1$) regimes in terms of dimensional quantities. Recall that $\kappa$ is defined as $\kappa=B\mathcal{T}/(a^{2}R\mu)$, where $\mathcal{T}$ is the timescale for the head motion, $a$ and $R$ are respectively the small and large radii defining the canals, and $B$ is the cupula’s bending stiffness. We see that the soft cupula regime may be attained for fast movements (small $\mathcal{T}$), large canals ($a,R$ large) and of course soft cupulas in absolute terms (small $B$). The converse holds for the stiff cupula regime. Inserting parameter values for $a$, $R$, $B$ and $\mu$ into the transition value predicted by our model, $\kappa_{c}=16I_{4}\alpha(b)$, we obtain a critical value
$\mathcal{T}=\frac{a^{2}R\mu\kappa_{c}}{B}=\frac{12(1-\nu^{2})\,a^{2}R\mu\,\kappa_{c}}{E\,t_{h}^{3}}\sim\frac{a^{2}R\mu}{Et_{h}^{3}},$ (58)
which may be interpreted as a critical frequency of rotation below which the system responds to angular velocity, and above which the system responds to angular acceleration. Inserting typical values for a human adult, we compute a transition frequency of 0.27 Hz. Interestingly, human experiments with controlled oscillation frequencies have reported a maximum motion sickness when the frequency is around 0.2 Hz (Golding et al., 2001). Our analysis suggests an intriguing possible explanation for this maximal sickness at intermediate frequencies: intermediate frequencies correspond to motion for which the response of the cupular system follows neither the angular acceleration nor velocity. The “neural mismatch” hypothesis predicts that motion sickness is induced in situations where there is a disagreement between visual or vestibular cues and the information anticipated by the nervous system (Benson, 1990). Since the vestibular system may be expected to provide information that matches neither the true acceleration or the true velocity around the transition point we suggest that it may be the transition point between small and large $\kappa$ that causes motion sickness.
Evidence suggests that susceptibility to motion sickness peaks at a higher frequency for animals smaller than humans Golding & Gresty (2016); Javid & Naylor (1999). Given the preceding discussion, this is a little surprising: if we assume that the size scales of the vestibular system scale in proportion then (58) shows that the critical frequency should remain the same. If the preceding hypothesis is correct, it would suggest that either the material parameters of the cupula change or that some non-trivial allometric scaling of the dimensions of the canal must occur. (We are unaware of any data on the allometric scaling.)
In §3.2 we also mentioned that in the transition region (when $\kappa$ is nether large nor small) the response develops an overshoot at the end of the manoeuvre: the deformation, and hence a sensing signal persists after the motion has concluded. This feature seems undesirable (the “neural mismatch” hypothesis would predict a high likelihood of experiencing motion sickness, as the vestibular input will disagree with the visual input) but also in line with everyday experience of dizziness.
From an engineering perspective, the insights gained from this study could inform the design of biomimetic sensors, prosthetics and systems, for example the MEMS prototype from Raoufi et al. (2019). For instance, understanding the interplay between fluid dynamics and flexible structures in the semicircular canals could inspire the development of flow sensors or inertial measurement devices that mimic the sensitivity and robustness of the vestibular system, for instance, biologically-inspired inertial navigation systems. Additionally, the analytical framework developed here could be extended to other biological systems involving thin fluid-filled structures, such as the cochlea or cardiovascular vessels.
7.2 Limitations and future directions
Our model has been based on a number of approximations that have facilitated the analysis that we have presented here. Of these, perhaps the most important is our use of a thin plate model of the cupula — this is mathematically convenient but is likely to be an over-simplification of the true behaviour of the cupula. In particular, its validity would require the cupula to be slender; concrete measurements of the cupula’s aspect ratio are not available, meaning our thin-plate model of the cupula might not be appropriate in all cases. As an alternative strategy, we note that the strains are small, and so a solution of the Navier equations (2) that does not assume the thickness of the cupula to be small might be required. At the same time, we note that our existing analysis assumes that the cupula’s thickness is constant, but photographic evidence suggests this is likely not the case (Rabbitt et al., 2004).In particular, the cupula seems to be thinner in the centre and thicker towards the edge. Modelling this effect, via a plate equation with a spatially dependent thickness, might be enough to warrant further work with the plate model. Whilst this will influence the shape of the deflection profile of the cupula, it is unlikely to give rise to new phenomena.
The model we have presented allows for the cupula to have some pre-tension, $T$, although this has not been observed experimentally. This lack of direct observation is in part because of the great difficulty in accessing this delicate tissue. (Pre-tension is often present in thin biological structures due to residual stresses from development or mechanical interactions with surrounding tissues (Goriely, 2017).) Experiments that determine the presence of pre-tension in the cupula (and its magnitude) could prove useful in understanding the role of parameters like the inverse bendability, $b$.
Another simplification in our model is that that the cupula is uniformly clamped to the crista and canal walls. However, previous results based on this assumption Selva et al. (2009) required a very small Young’s modulus, close to $5\mathrm{~{}Pa}$, to match experimentally observed deformations Selva et al. (2009). This is an extremely low value, perhaps indicating the softest material in the human body, and is unrealistic when compared to other “soft” biological tissues (Goriely, 2017). We suggest that this anomalous stiffness of the cupula may be a result of the clamped boundary conditions on all sides of the cupula, as used here and as usual in the vestibular literature (Rabbitt & Damiano, 1992), may be incorrect; typical anatomical drawings suggest that the cupula is only clamped on a part of its boundary, and is free to move on other regions of the boundary. This would increase the apparent flexibility of the cupula, creating similar system behaviour with without requiring an unusually small Young’s modulus.
Our model may also allow for other phenomena within the vestibular system to be investigated. An interesting avenue using the techniques developed here is the light cupula phenomenon (see Lee et al., 2024, for a review of the concept), as well as related concepts such as the buoyancy hypothesis to explain balance loss after alcohol intake Brandt (1991). These essentially state that when alcohol is consumed, ethanol diffuses faster into the cupula than the surrounding endolymph, changing their density ratio (which under normal functioning is very close to one, so that the cupula is neutrally buoyant). As ethanol is less dense than water, the cupula would then become negatively buoyant, deforming differently than ordinarily and sending incorrect signals to the nervous system. To account for this in our model, we would need to include a buoyancy term so that the cupula can float or sink through the endolymph.
[…]
[Acknowledgements]We acknowledge fruitful conversations with Dr Miguel Vaca, Dr Eduardo Martin Sanz, Prof. Michael Gresty and Prof. Adolfo Bronstein. We are further grateful to Prof. Sarah L. Waters and Prof. Ian M. Griffiths for their invaluable feedback on an early version of the project. For the purpose of Open Access, the authors will apply a CC BY public copyright license to any Author Accepted Manuscript version arising from this submission.
[…]
[Funding]JCV is funded by a St. John’s College scholarship.
[…]
[Declaration of interests]The authors report no conflict of interest.
[…]
[Data availability statement]The data that support the findings of this study will be made openly available in the Oxford Research Archive.
[…]
[Author ORCIDs]Javier Chico-Vazquez, https://orcid.org/0009-0008-7536-6978.
Dominic Vella, https://orcid.org/$0$-0003-1341-8863
Derek Moulton, https://orcid.org/$0$-0003-3597-7973
Appendix A Details for numerical solutions
The governing equations from Section 2 were solved in COMSOL for different values of the Young’s modulus. The equations were solved on a moving grid, without neglecting the geometric nonlinearity and including the nonlinear terms in the Navier–Stokes equations. As mentioned previously, the cupula is modelled as a full three-dimensional solid, without assuming it is a thin structure and with a finite thickness, $t_{h}$.
Figure 12: Numerically obtained deformation of the solid material at $\hat{t}=0.25$ s, normalized by the maximum deformation. Left panel: deformation along the direction perpendicular to the canal centre-line (in the model’s coordinate system this is along $r$). We plot the membrane limit, $b\to 0$ (dashed black line) and the plate limit $b\to\infty$ (dotted line) for comparison. Middle panel: deformation in the direction parallel to the centreline (i.e. along $s$). Note the small change in deformation along $s$, suggesting a plate equation model that only tracks the deformation of the centre of the cupula is appropriate. Right panel: deformation at the centre of the cupula plotted as a function of the Young’s modulus.
In Figure 12 we plot the deformation of the cupula, by showing (left panel) the deformation of the cupula in the direction normal to the flow and (middle panel) the deformation of the cupula in the direction along the flow. In the first case, we observe that the deformation is large in the centre of the structure and zero on the edges, as expected. However, the structure does not seem to be clamped as might have been expected — the edge exhibits a finite gradient. This disagreement might be an effect of the finite thickness of the cupula used in the numerical simulations, or alternatively, might occur because of numerical artifacts and lack of resolution near the edge. For completeness, we have also included the analytical solution of the plate case ($b\to\infty$) with pinned boundary conditions. This solution does not match the numerical solution either, even if it has a finite gradient at the boundary. Following physical intuition and previous work (Rabbitt & Damiano, 1992), we have opted to present the clamped solution in the main text.
In the left panel fig. A we plot the normalized deformation, and we observe all curves collapse. This indicates the deformation regime is linear and motivates the use of a linear plate equation in the theoretical model of the main text. In the second plot, we see that the deformation only varies by around 2% in the direction parallel to the centreline. Again, this is consistent with the use of a linear plate equation where only the deformation of the centre plane of the cupula is tracked. (The forcing used to generate Figure 12 is that given by (34).) Finally, we note that the magnitude of the deformation is inversely proportional to the stiffness $E$, as can be seen in Figure 12 (right panel).
Appendix B Numerical procedure for integro-differential equations
When the Stokes number of the flow is no longer negligible, the deformation of the cupula satisfies (43), which is a Volterra integro-integral equation(Polianin & Manzhirov, 1998), and so may be solved numerically using the trapezoidal method. For small values of the Stokes number this solution procedure requires an increasingly fine temporal discretization, making the problem computationally intensive. Therefore, an alternative numerical scheme is required.
Integral equations with exponential Kernels can be transformed into systems of ODEs by introducing additional variables (Wazwaz, 2011); although the kernel $\mathcal{K}(x,\mathrm{St})$ in (43) is not strictly exponential, it may be seen as a linear combination of exponential Kernels. To this end, we may define a sequence of auxiliary variables
$z_{n}(t)=\frac{1}{\mathrm{St}\lambda_{n}^{2}}\int_{0}^{t}\left(f(t)+\frac{\kappa}{2\pi}\Delta p_{0}\right)e^{-\lambda_{n}^{2}/\mathrm{St}(t-\tau)}~{}\mathrm{d}\tau,\quad n=0,\dots,N-1.$ (59)
Upon truncation of the infinite series (43) reads
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-2\sum_{n=0}^{N-1}z_{n}.$ (60)
Considering $\frac{\mathrm{d}z_{n}}{\mathrm{d}t}$ and differentiating under the integral sign we find that
$\frac{\mathrm{d}z_{n}}{\mathrm{d}t}=\frac{1}{\mathrm{St}\lambda_{n}^{2}}\left(f(t)+\frac{\kappa}{2\pi}\Delta p_{0}\right)-\frac{\lambda_{n}^{2}}{\mathrm{St}}z_{n}.$ (61)
Therefore we have a system of $N+1$ differential equations for the $N+1$ unknowns, which may be solved efficiently for any value of the Stokes number.
Appendix C Laplace transform approach for finite Stokes number
The general equation determining the shape of the cupular deflection is (40), which we Laplace transform in time to obtain:
$\int_{0}^{a_{0}}r\sigma\tilde{\eta}_{0}\mathrm{d}r$ $=-2a(s)^{2}\frac{1}{\mathrm{St}}\sum_{n=0}^{\infty}\frac{1}{\lambda_{n}^{2}}\left(\tilde{f}(\sigma)+\kappa\frac{\partial\tilde{p}_{0}}{\partial s}\right)\frac{1}{\sigma+\lambda_{n}^{2}/(a^{2}\mathrm{St})},$ (62a)
$\nabla^{4}$ $\tilde{\eta}_{0}-b^{-1}\nabla^{2}\tilde{\eta}_{0}+\varrho\left(\mathrm{St}\sigma^{2}\tilde{\eta}_{0}+\tilde{f}(\sigma)\right)=\Delta\tilde{p}_{0}.$ (62b)
(Here the convolution theorem has been used to compute the transform of the convolution integral.) We may now isolate the pressure gradient in the first equation as it may be factored out of the sum
$\tilde{f}(\sigma)+\kappa\frac{\partial\tilde{p}_{0}}{\partial s}=-\frac{\mathrm{St}\sigma}{2a^{2}}\int_{0}^{a_{0}}r\tilde{\eta}_{0}~{}\mathrm{d}r\frac{1}{\sum_{n=0}^{\infty}\lambda_{n}^{-2}\left(\sigma+\lambda_{n}^{2}/(a^{2}\mathrm{St})\right)^{-1}}.$ (63)
Hence, after integrating along the length of the duct,
$2\pi\tilde{f}(\sigma)+\kappa\Delta\tilde{p}_{0}=-\frac{\sigma}{2}\int_{0}^{a_{0}}r\tilde{\eta}_{0}\mathrm{d}r\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}\sum_{n=0}^{\infty}\lambda_{n}^{-2}\left[a(s)^{2}\mathrm{St}\sigma+\lambda_{n}^{2}\right]^{-1}}.$ (64)
Substituting the pressure jump using the transformed plate equation leads to a single equation for $\tilde{\eta}_{0}$,
$2\pi\tilde{f}(\sigma)+\kappa\left[\nabla^{4}\tilde{\eta}_{0}-b^{-1}\nabla^{2}\tilde{\eta}_{0}+\varrho\left(\mathrm{St}\sigma^{2}\tilde{\eta}_{0}+\tilde{f}(\sigma)\right)\right]=-\frac{\sigma}{2}\int_{0}^{a_{0}}r\tilde{\eta}_{0}\mathrm{d}r\frac{1}{\tilde{\mathcal{K}}(\sigma;\mathrm{St})},$ (65)
where the transformed Kernel is
$\tilde{\mathcal{K}}(\sigma;\mathrm{St})=\left(\int_{0}^{2\pi}\frac{\mathrm{d}s}{a(s)^{4}\sum_{n=0}^{\infty}\lambda_{n}^{-2}\left[a(s)^{2}\mathrm{St}\sigma+\lambda_{n}^{2}\right]^{-1}}\right)^{-1}.$ (66)
Multiplication by the transformed Kernel, followed by the inversion of the transform and the application of the convolution theorem leads to
$\int_{0}^{a_{0}}r\frac{\partial\eta_{0}}{\partial t}\mathrm{d}r=-4\pi\int_{0}^{t}\left(f(\tau)+\frac{\kappa}{2\pi}\left[\nabla^{4}\eta_{0}-b^{-1}\nabla^{2}\eta_{0}+\varrho\left(\mathrm{St}\frac{\partial^{2}\eta_{0}}{\partial t^{2}}+f(\tau)\right)\right]\right)\mathcal{K}(t-\tau)\mathrm{d}\tau,$ (67)
where $\mathcal{K}(t;\mathrm{St})=\mathcal{L}^{-1}[\tilde{\mathcal{K}}(\sigma;\mathrm{St})]$ is the Kernel. When the cupula’s inertia is negligible ($\varrho\ll 1$), we write the deflection as $\eta_{0}(r,t)=\bar{\eta}(r)\Delta p_{0}(t)$, where $\nabla^{4}\bar{\eta}(r)-b^{-1}\nabla^{2}\bar{\eta}(r)=1$ and
$\alpha(b)\frac{\mathrm{d}\Delta p_{0}}{\mathrm{d}t}=-4\pi\int_{0}^{t}\left(f(\tau)+\frac{\kappa}{2\pi}\Delta p_{0}(\tau)\right)\mathcal{K}(t-\tau;\mathrm{St})\mathrm{d}\tau.$ (68)
This calculation is crucial because it systematically reduces the governing equation for the cupular deflection into a solvable integral equation by leveraging the Laplace transform. By transforming the original time-dependent equations, the problem is converted into an algebraic form where the pressure gradient can be explicitly isolated, and integrated in space to obtain the pressure jump. Furthermore, inverting the transform and applying the convolution theorem ultimately yields an explicit time-domain equation governing the evolution of $\eta_{0}$. This final equation is particularly useful, as it expresses the cupular deflection in terms of a convolution integral. This approach allows for the computation of the solution in arbitrary domains, linking the problem to that solved in a simpler domain via the transformed Kernel.